OpenStudy (laweejay):

Help! The coefficients of static and kinetic friction are μs = 0.73 and μk = 0.62, respectively.The acceleration of gravity is 9.8 m/s2 . What is the frictional force acting on the 49 kg mass? The angle of incline is 34 degrees. Answer in units of N

OpenStudy (vincent-lyon.fr):

Please, post a sketch.

OpenStudy (laweejay):



OpenStudy (shamim):

Weight=mg=49*9.8N Acting along just downward

OpenStudy (shamim):

Component of weight along inclined surface is= 49*9.8*sin34

OpenStudy (shamim):

Component of weight perpendicular to the inclined plane is=49*9.8*cos34

OpenStudy (shamim):

Static frictional force=coefficient of static friction*R=0.73*49*9.8*cos34=?

OpenStudy (shamim):

Feel free to ask for getting better help!!!

OpenStudy (laweejay):

Thanks!

OpenStudy (shamim):

U r most welcome!!!

OpenStudy (shamim):

Meda???!!!!!